OpenStudy (anonymous):

What is the ratio of x to y? 2/3 = x/3x+y A. -3/2 B. -2/3 C. 2/3 D. 3/2

OpenStudy (dayakar):

\[\frac{ 2 }{ 3} =\frac{ x }{ 3x+y }\] do the cross multiplication 2(3x+y) = 3x 6x +2y = 3x 6x-3x = -2y 3x = -2y find remaining part

OpenStudy (anonymous):

okay thank you

OpenStudy (dayakar):

what is the answer?

OpenStudy (anonymous):

idk

OpenStudy (dayakar):

3x =-2y multiply both sides with 1/3y \[\frac{ 1 }{ 3y }\times 3x = \frac{ 1 }{ 3y } \times (-2y)\] x/y = -2/3

OpenStudy (anonymous):

Oh okay thank you im sorry

OpenStudy (dayakar):

do u understand go through the solution step by step

OpenStudy (anonymous):

Yes okay i will thank you for helping me understand i appreciate your help

OpenStudy (dayakar):

:)